Evaluate the integral. IP In(r) dr
Answer #1

983 ln (8) dr į fordr = F(b) – F(a) ; F(Y) = \ tus or F(x) = J q? h (r) dr Integration by parts Ju.v=uu-fur let u = locor) ;